OpenStudy (anonymous):

find non zero matrices A,B and C such that AC=BC and A is not equal to B find non zero matrices A,B and C such that AC=BC and A is not equal to B @Mathematics

OpenStudy (perl):

hmmm

OpenStudy (perl):

so we know at least that C cannot be invertible because if C was invertible we could do A*C*C^-1 = B*C*C^-1 and then A = B

OpenStudy (anonymous):

please explain by taking any example

OpenStudy (anonymous):

try \[C= \begin{matrix} 1& 1 \\ 1 & 1 \\ \end{matrix} \]

OpenStudy (zarkon):

didn't I already answer this question

OpenStudy (anonymous):

oh i didn't see it

OpenStudy (anonymous):

yes u have already told me zarkon but u do not tell me how A and B matrices are taken with what limitations and i want to clear that.

OpenStudy (perl):

determinant of C is zero